You dump 255 g of unknown metal (at 125 C) into a 955 g vat of water at 20.0 C, and the

final temperature of both is 35.7 C. What is the specific heat of the unknown? Assume NO water I lost...
Answer #1

Here ,

let the specific heat of unknown is S

for the heat

heat lost by metal = heat lost by water

955 * 4.186 * (35.7 - 20) = 255 * S * (125 - 35.7)

solving for S

S = 2.76 J/(gm. degree C) = 2760 J/(kg.degree C)

the specific heat of unknown is 2760 J/(kg.degree C)